Mr.carlson drove a distance of 1000 miles to another city. He returned by plane. The eate of the plane was 10 times the rate at which he drove. He spent 22 hours on the trip. Find the rate at which he drove and the rate of the plane
Answer by nerdybill(7384)

Mr.carlson drove a distance of 1000 miles to another city. He returned by plane. The rate of the plane was 10 times the rate at which he drove. He spent 22 hours on the trip. Find the rate at which he drove and the rate of the plane
.
Let x = speed at which Mr. Carlson drove
then
10x = speed of the plane
.
At this point we use the "distance formula":
D=RT
where
D is distance
R is rate or speed
T is time
.
In our case, we're going to solve for T:
D=RT
T = D/R
.
"time taken to drive" + "time taken to fly" = "total time of travel"
.
1000/x + 1000/10x = 22
multiplying both sides by 10x we get:
10000 + 1000 = 220x
11000 = 220x
50 mph = x (speed at which he drove)
.
10x = 10*50 = 500 mph (speed of plane)